Express the following comparative income statements in common-size percents and assess whether this company’s situation has...

  1. Express the following comparative income statements in common-size percents and assess whether this company’s situation has improved in the current year.

Comparative Statements (in thousands of $)

          For the Years Ended December 31                                                 Current Year                       Prior Year

                Sales                                                                                                      $800                                       $500

                Total Expenses                                                                                 $560                                       $400

                Net Income                                                                                        $240                                       $100

  1. Using the following account balances and other information from Company “A” accounting records for year-end December 31, 2019. Prepare a multi-step income statement.
Answer #1

Income Statement

Particular Current Year Percentage Prior Year Percentage
Sales $800 100% $500 100%
Less: Expenses $560 $560/$800*100= 70% $400 $400/$500*100=80%
Net Income (Sales-Expenses) $240 $240/$800*100=30% $100 $100/$500*100= 20%

Assessment= One can see that profit percentage have been improved since last year.So, company's situation have beem improved.